OpenStudy (anonymous):

Suppose that you must order concrete for a sidewalk 51 ft by 2 ft to a depth of 7 in. How much concrete is required? (Answer to the nearest 1/2 cubic yard.)

OpenStudy (anonymous):

If you convert each measurement to inches, knowing there 12 inches in each ft. You would have 602in by 24in by 7in. This would be like a volume problem. 602*24*7 = 101136 cubic inches. There are 36 inches in a yard, so a cubic yard would be 36*36*36 = 46656 so 101136/46656 = 2.0 cubic yards